# SOUL: Unlocking the Power of Second-Order Optimization for LLM Unlearning

### v3

@@ 5.1 Experiment setups
Datasets, tasks and models.
Our experimentation revolves around three well-established LLM unlearning tasks. (1) TOFU: This task focuses on fictitious unlearning Maini et al. (2024), involving a dataset of fictitious author profiles for finetuning, and a subset of these profiles constitutes the forget set. (2) Copyrighted information removal: This task evaluates the effectiveness of unlearning methods in reducing potential copyright infringement Yao et al. (2023); Eldan and Russinovich (2023). (3) Model detoxification: This task aims to prevent LLMs from generating toxic content Yao et al. (2023); Ilharco et al. (2022); Zhang et al. (2023c) by employing unlearning approaches. In the TOFU task, we form a forget set by selecting a 10% forget ratio, which includes 400 examples providing information about 20 authors, along with the remaining data points to form the retain set. In the task of removing copyrighted information, we extract 200 chunks from the Harry Potter book series dataset Eldan and Russinovich (2023), with each chunk containing up to 512 tokens, to create the forget set. For the model detoxification task, we include 200 negative samples from the PKU-SafeRLHF training set Ji et al. (2024) as the forget set. The C4 dataset Raffel et al. (2020) is used as the retain set for copyright removal and model detoxification tasks to ensure the preservation of model utility.
As for model configurations, We use the OPT-1.3B Zhang et al. (2022b), and LLaMA2-7b Touvron et al. (2023) as our base models. In the TOFU task, we specifically utilized the fine-tuned LLaMA2-7b-chat model. In the copyright removal task, we fine-tuned both the OPT-1.3B and LLaMA2-7b models on the Harry Potter book series dataset Eldan and Russinovich (2023) to simulate the generation of copyrighted information. In the detoxification task, the original models were used without any additional modification. Further information regarding model preparation and configuration can be found in Appendix A.1.
@@
Forget quality Acc. Rouge-L Acc. Rouge-L Acc. Rouge-L Acc. Rouge-L
Original 0.36 85.25% 0.9796 85.75% 0.9825 89.00% 0.9330 86.32% 0.8960
Input-based 0.30 79.50% 0.6536 77.50% 0.6651 64.00% 0.6480 77.78% 0.8205
FO-GA 0.14 66.25% 0.4110 63.25% 0.4504 42.00% 0.4400 76.92% 0.8170
FO-GradDiff 0.02 72.75% 0.5174 76.50% 0.6115 71.00% 0.7677 79.49% 0.8462
SO-GradDiff (Ours) 1.00 10.25% 0.0221 72.25% 0.5960 78.00% 0.8113 82.05% 0.8675
FO-PO 0.72 37.00% 0.0882 82.75% 0.9051 90.00% 0.9330 84.62% 0.8875
SO-PO (Ours) 0.92 28.75% 0.0761 82.75% 0.8137 90.00% 0.9380 86.32% 0.9046
Table 1: Overview of the fictitious unlearning performance using different LLM unlearning approaches under the TOFU fine-tuned LLaMA2-7B-chat model Maini et al. (2024). ‘Original’ refers to the original model without unlearning. ‘FO’ and ‘SO’ indicate the choice of the unlearning optimizer, either FO unlearning or SOUL. As illustrated in experiment setups, the algorithmic frameworks of LLM unlearning include GA, GradDiff, and PO. The proposed second-order LLM unlearning methods correspond to SO-GradDiff and SO-PO. The symbol denotes metrics where lower values indicate better unlearning performance, while symbolizes metrics where higher values are preferable, reflecting better retention of model utility. The ‘Unlearning Efficacy’ category measures the model’s success in removing targeted information, whereas ‘Utility’ gauges the model’s retained functionality post-unlearning. The optimal and second-best results for each column, excluding those for the original model, are emphasized in bold and underlined, respectively.
LLM unlearning methods and implementations.
We will assess the effectiveness of our proposed second-order unlearning approach by comparing it with a series of state-of-the-art (SOTA) LLM unlearning techniques as outlined below. Gradient ascent (GA): This serves as a specialization of GradDiff (4) by setting its regularization parameter $\lambda=0$. GradDiff (4) and PO (6) are executed via regularized optimization, employing either FO (first-order) or SO (second-order) optimizers. In the implementation of PO, we choose a reject-based answer as the target response $y_{\mathrm{f}}$ to steer the model away from unwanted responses. Table A1 in Appendix A.2 provides a summary of the reject-based answers utilized across various unlearning tasks. In addition to the aforementioned finetuning-based unlearning methods, we also explore an input prompt-enabled unlearning approach proposed by Thaker et al. (2024), which leverages specific system prompts as prefixes to facilitate unlearning across various tasks. Further details on these system prompts are provided in Table A2 of Appendix A.3. AdamW Loshchilov and Hutter (2017) is used as the FO optimizer, and Sophia Liu et al. (2023a) is utilized as the SO optimizer in our proposed SOUL framework presented in Algorithm 1. For detailed information on the hyperparameters used in our implementations, please refer to Appendix A.4.
@@
To evaluate the effectiveness of fictitious unlearning in the TOFU task, we measure the distinguishability of statistical measures between the forget and retain sets using LLM-generated truthful ratios, as defined in the original TOFU benchmark Maini et al. (2024). This assessment is conducted via the Kolmogorov-Smirnov (KS) test. We utilize $1-$ $p$-value obtained from the KS test as the Forget Quality to assess unlearning effectiveness. In the experimentation, a high forget quality represents successful unlearning, indicating an increased distributional divergence between the forget and retain sets. This aligns with the objectives of membership inference attacks (MIA) Song et al. (2019); Hayes et al. (2024) in unlearning, which aim to determine whether individual data points in the forget set were part of the training dataset. Additionally, we assess the performance of the LLM after unlearning (referred to as the unlearned model) by computing the Rouge-L recall against the ground truth and measuring the accuracy of the generated text. This involves comparing the cosine similarity of semantic embeddings from Sentence-BERT Reimers and Gurevych (2019) with both the ground truth and alternative incorrect responses in the TOFU dataset. Correctness is determined when the semantic embedding of the generated text is closest to the ground truth. We apply the same accuracy and Rouge-L recall metrics to evaluate utility preservation on sets related to retained information, real authors, and world facts.
In the copyright removal task, we randomly truncate 300 excerpts from the original Harry Potter dataset to the first $k$ tokens and evaluate them using BLEU and Rouge-L recall for prompt lengths of 100 and 300 tokens, with text completion instructions found in Appendix A.5. In the model detoxification task, toxicity is assessed using real toxic prompts Gehman et al. (2020) and the PKU-SafeRLHF test set Ji et al. (2024), assigning toxicity scores with Toxic-BERT Hanu and Unitary team (2020). For both the copyright removal and detoxification tasks, utility preservation is assessed using the LM Evaluation Harness Gao et al. (2023) to compute perplexity (PPL) on the Wikitext Merity et al. (2016) and mean zero-shot accuracy across tasks, detailed in Appendix A.6. Additional evaluations include TruthfulQA Lin et al. (2021).
@@ 5.2 Results on fictitious unlearning in TOFU
In Table 1, we showcase the unlearning effectiveness and the preserved model utility following the application of various LLM unlearning methods to the TOFU fine-tuned LLM Maini et al. (2024), with a focus on comparing FO (first-order) unlearning with the proposed SO unlearning, SOUL. As we can see, SOUL-based methods significantly outperform their FO counterparts (FO-GradDiff vs. SO-GradDiff, and FO-PO vs. SO-PO) in the efficacy measurements of LLM unlearning. This is evident from the enhanced forget quality, accuracy, and Rouge-L scores on the forget set. Moreover, SOUL-based methods effectively preserve the model’s utility post-unlearning. This is evident from their competitive utility performance compared to FO-GradDiff or FO-PO, as well as the improvement over FO-GA and the input prompt-oriented unlearning method Thaker et al. (2024). Among the studied unlearning methods, SO-PO (second-order preference optimization) appears to strike a graceful balance between unlearning effectiveness and utility preservation.
@@
Table 2: Example of generated texts from different unlearned models in the TOFU dataset. Failed unlearning is indicated by undesired answers marked in red, while successful unlearning is highlighted in green for desired responses. More examples are provided in Appendix A.7.
@@ 5.3 Results on copyright removal
Table 3 presents the unlearning efficacy and model utility of the proposed SO unlearning methods and baselines in the task of ‘Who’s Harry Potter’ copyrighted information removal across two LLMs fine-tuned on the Harry Potter book series dataset Eldan and Russinovich (2023). Consistent with our observations in the TOFU task, SOUL substantially improves the unlearning efficacy. For example, the comparison between FO-GradDiff and SO-GradDiff shows a notable decrease in BLEU score (by 0.21) at a prompt length of 300 in the LLaMA2-7B model. This decrease suggests that the generated texts deviate further from the original book’s content. Furthermore, the enhancements observed in both perplexity (PPL) and zero-shot accuracy with SOUL over FO unlearning highlight a superior balance between forget efficacy and utility preservation. Similar to the TOFU task, the GA method struggles to balance forget efficacy with utility preservation. Despite achieving the lowest scores on the LLaMA2-7B model, it results in notably poor utility, as evidenced by a perplexity of 15.66, substantially higher than other methods. Table A5 in Appendix A.7 showcases visualization examples, further demonstrating the enhanced performance of SOUL.
@@
BLEU ROUGEL BLEU ROUGEL
@@
5.4 Results on LLM detoxification
In Table 4, we demonstrate that the proposed SO unlearning methods effectively reduce the toxicity score on both the Real Toxicity Prompts and PKU-SafeRLHF datasets while maintaining or even improving utility. For instance, in the LLaMA2-7B model, SO-PO achieved a clear reduction in the toxic score on the PKU-SafeRLHF dataset and showed enhanced performance in zero-shot accuracy compared to FO-PO. This indicates improved unlearning efficacy of SOUL without sacrificing model utility. In addition, Table A6 includes visualizations that exemplify the outputs after the application of unlearning to the LLaMA2-7B models. These visualizations further corroborate that SO optimizers improve unlearning efficacy, particularly highlighting that SO-PO achieves the most effective unlearning performance.
Method Forget efficacy Utility
Real Toxicity Prompts PKU-SafeRLHF PPL Zero-shot Acc. TruthfulQA
Toxic Score Toxic Score
OPT-1.3B
Original 0.0807 0.1118 16.49 48.16% 0.2411
FO-GradDiff 0.0748 0.0673 30.87 41.16% 0.2362
SO-GradDiff (Ours) 0.0561 0.0618 28.77 40.34% 0.2240
FO-PO 0.0404 0.0253 18.26 46.25% 0.2852
SO-PO (Ours) 0.0335 0.0165 17.97 48.60% 0.2742
LLaMA2-7B
Original 0.0710 0.1027 8.79 62.08% 0.2521
FO-GradDiff 0.0708 0.0989 8.77 61.38% 0.2534
SO-GradDiff (Ours) 0.0722 0.0987 8.79 61.32% 0.2534
FO-PO 0.0626 0.0790 8.78 61.92% 0.2632
SO-PO (Ours) 0.0528 0.0443 8.87 62.80% 0.2656
Table 4: Performance comparison between SOUL and its FO counterparts in the task of model detoxification, following the format of Table 3.
6 Conclusions
@@ 8 Acknowledgement
Appendix A Additional Experimental Details and Results
A.1 Model configurations.
We selected the OPT-1.3B Zhang et al. (2022a), and LLaMA2-7b Touvron et al. (2023) as foundational models for our study. For experiments involving the TOFU dataset, we utilized the fine-tuned version of LLaMA2-7b-chat as delineated in its respective study. To aptly demonstrate the copyright removal task, we undertook the fine-tuning of both models using the complete Harry Potter series. The fine-tuning procedure for the OPT-1.3B model involved a learning rate of $5\times 10^{-5}$ and a batch size of 2. Conversely, for LLaMA2-7b, we applied Low-Rank Adaptation (LoRA) fine-tuning with a learning rate of $1\times 10^{-4}$ and the same batch size. AdamW served as the optimizer for preparing these models. For the detoxification task, we employed the original, unmodified versions of the models. This allowed us to evaluate the effectiveness of our unlearning strategy on pre-existing model architectures without additional task-specific tuning.
A.2 Reject-based answers in PO.
Table A1 shows the reject-based answers we designed in the preference optimization method:

### v4

@@ 5.1 Experiment setups
Unlearning tasks and models.
Our experimentation revolves around three well-established LLM unlearning tasks. (1) TOFU: This task focuses on fictitious unlearning Maini et al. (2024), involving a dataset of fictitious author profiles for finetuning, and a subset of these profiles constitutes the forget set (with 10% forget ratio). (2) Copyrighted information removal: This task evaluates the effectiveness of unlearning methods in reducing potential copyright infringement Eldan and Russinovich (2023). (3) Model detoxification: This task aims to prevent LLMs from generating toxic content Yao et al. (2023); Ilharco et al. (2022); Zhang et al. (2023c) by employing unlearning approaches. To achieve these unlearning tasks, we use the OPT-1.3B Zhang et al. (2022b) and LLaMA2-7b Touvron et al. (2023) as our base models. We refer readers to Appendix B.1 for more details on the tasks, datasets, and model configurations.
LLM unlearning methods.
We will assess the effectiveness of our proposed second-order unlearning approach by comparing it with a series of state-of-the-art (SOTA) LLM unlearning techniques. As illustrated in Sec. 3, we consider GradDiff, PO, and NPO, executed via regularized optimization and employing either FO (first-order) optimization or SOUL. We also consider Gradient ascent (GA), which serves as a specialization of GradDiff (4) by setting its regularization parameter $\lambda=0$. In addition to the aforementioned finetuning-based unlearning methods, we also explore an input prompt-enabled unlearning approach proposed by Thaker et al. (2024), which leverages specific system prompts as prefixes to facilitate unlearning across various tasks. We refer readers to Appendix B.2 for more implementation details.
Tasks Efficacy/Utility Metrics
TOFU Unlearning efficacy Forget quality $\uparrow$
Accuracy on forget set $\downarrow$
Rouge-L on forget set $\downarrow$
Membership inference attack $\downarrow$
Utility Accuracy on retain set $\uparrow$
Rouge-L on retain set $\uparrow$
Accuracy on real author set $\uparrow$
Rouge-L on real author set $\uparrow$
Accuracy on world facts set $\uparrow$
Rouge-L on world facts set $\uparrow$
Copyrighted information removal Unlearning efficacy BLEU on Harry Potter completion $\downarrow$
Rouge-L on Harry Potter completion $\downarrow$
Utility Perplexity on Wikitext $\downarrow$
Zero-shot Accuracy on benchmarks
$\uparrow$
Zero-shot Accuracy on TruthfulQA $\uparrow$
Detoxification
Unlearning
efficacy
Toxic score $\downarrow$
Utility Perplexity on Wikitext $\downarrow$
Zero-shot Accuracy on benchmarks
$\uparrow$
Zero-shot Accuracy on TruthfulQA $\uparrow$
Table 1: Summary of unlearning effectiveness metrics and model utility metrics used for different LLM unlearning tasks. The or indicates whether a lower or higher value is desired for better performance, respectively.
@@
Forget quality Acc. Rouge-L MIA Acc. Rouge-L Acc. Rouge-L Acc. Rouge-L
Original 0.36 85.25% 0.9796 0.7894 85.75% 0.9825 89.00% 0.9330 86.32% 0.8960
Input-based 0.30 79.50% 0.6536 0.7894 77.50% 0.6651 64.00% 0.6480 77.78% 0.8205
FO-GA 0.14 66.25% 0.4110 0.7754 63.25% 0.4504 42.00% 0.4400 76.92% 0.8170
FO-GradDiff 0.02 72.75% 0.5174 0.7627 76.50% 0.6115 71.00% 0.7677 79.49% 0.8462
SO-GradDiff (Ours) 1.00 10.25% 0.0221 0.2156 72.25% 0.5960 78.00% 0.8113 82.05% 0.8675
FO-PO 0.72 37.00% 0.0882 0.7911 82.75% 0.9051 90.00% 0.9330 84.62% 0.8875
SO-PO (Ours) 0.92 28.75% 0.0761 0.7877 82.75% 0.8137 90.00% 0.9380 86.32% 0.9046
FO-NPO 1.00 16.00% 0.0458 0.3062 80.75% 0.8426 85.00% 0.9110 82.91% 0.8803
SO-NPO (ours) 1.00 16.00% 0.0291 0.2274 81.25% 0.8314 89.00% 0.9283 85.47% 0.8917
Table 2: Overview of the fictitious unlearning performance using different LLM unlearning approaches under the TOFU fine-tuned LLaMA2-7B-chat model Maini et al. (2024). ‘Original’ refers to the original model without unlearning. ‘FO’ and ‘SO’ indicate the choice of the unlearning optimizer, either FO unlearning or SOUL. As illustrated in experiment setups, the algorithmic frameworks of LLM unlearning include GA, GradDiff, PO, and NPO. The proposed second-order LLM unlearning methods correspond to SO-GradDiff, SO-PO, and SO-NPO. The symbol denotes metrics where lower values indicate better unlearning performance, while symbolizes metrics where higher values are preferable, reflecting better retention of model utility. The ‘Unlearning Efficacy’ category measures the model’s success in removing targeted information, whereas ‘Utility’ gauges the model’s retained functionality post-unlearning. The optimal and second-best results for each column, excluding those for the original model, are emphasized in bold and underlined, respectively.
@@
Table 1 summarizes the unlearning performance metrics, covering both unlearning effectiveness and preserved model utility across different LLM unlearning tasks. See more details on these metrics in Appendix B.3. We specify two unlearning effectiveness metrics, forget quality and membership inference attack (MIA), for the fictitious unlearning on TOFU, as their definitions were not covered in the original TOFU benchmark. First, forget quality characterizes the distinguishability of statistical measures between the forget and retain sets using LLM-generated truthful ratios. This assessment is conducted via the Kolmogorov-Smirnov (KS) test. We use $1-$ p-value from the KS test as the forget quality to assess unlearning effectiveness. A high forget quality represents better unlearning, indicating an increased distributional divergence between forget and retain sets. Second, MIA is achieved through the Min-kProbability method Shi et al. (2023). This method determines whether a specific piece of text was part of an LLM’s training dataset. For our evaluation, we measure the Area Under the Curve (AUC) of the Min-k%-based MIA detector to identify whether the forgotten data was originally included in the training set. A well-unlearned model should achieve a lower AUC, indicating improved effectiveness by not detecting forgotten data as part of the training set. Regarding utility, we did not consider more complex evaluations such as instruction-following ability. This is because the primary models are pre-trained, not adapted using RLHF Achiam et al. (2023).
@@ 5.2 Results on fictitious unlearning in TOFU
In Table 2, we showcase the unlearning effectiveness and the preserved model utility following the application of various LLM unlearning methods to the TOFU fine-tuned LLM Maini et al. (2024), with a focus on comparing FO (first-order) unlearning with the proposed SO unlearning, SOUL. As we can see, SOUL-based methods consistently outperform their FO counterparts (FO-GradDiff vs. SO-GradDiff, FO-PO vs. SO-PO, and FO-NPO vs. SO-NPO) in the efficacy measurements of LLM unlearning. This is evident from the improved forget quality, MIA, accuracy, and Rouge-L scores on the forget set. Moreover, SOUL-based methods effectively preserve the model’s utility post-unlearning. This is evident from their competitive utility performance compared to FO-GradDiff, FO-PO, and FO-NPO as well as the improvement over FO-GA and the input prompt-oriented unlearning method Thaker et al. (2024). Among the unlearning methods studied, SO-PO strikes a graceful balance between unlearning effectiveness and utility preservation. However, it falls short in achieving satisfactory results in MIA. This is because it does not explicitly reduce the Min-kprobability for the correct answer Shi et al. (2023), causing the data to still be recognized as a training example and leading to high MIA scores.
@@
FO-NPO
SO-NPO
Table 3: Example of generated texts from different unlearned models in the TOFU dataset. Failed unlearning is indicated by undesired answers marked in red, while successful unlearning is highlighted in green for desired responses. More examples are provided in Appendix B.4.
@@ 5.3 Results on copyright removal
Table 4 presents the unlearning efficacy and model utility of the proposed SO unlearning methods and baselines in the task of ‘Who’s Harry Potter’ copyrighted information removal across two LLMs fine-tuned on the Harry Potter book series dataset Eldan and Russinovich (2023). Consistent with our observations in the TOFU task, SOUL substantially improves the unlearning efficacy. For example, the comparison between FO-GradDiff and SO-GradDiff shows a notable decrease in BLEU score (by 0.21) at a prompt length of 300 in the LLaMA2-7B model. This decrease suggests that the generated texts deviate further from the original book’s content. Furthermore, the enhancements observed in both perplexity (PPL) and zero-shot accuracy with SOUL over FO unlearning highlight a superior balance between forget efficacy and utility preservation. Similar to the TOFU task, the GA method struggles to balance forget efficacy with utility preservation. Despite achieving the lowest scores on the LLaMA2-7B model, it results in notably poor utility, as evidenced by a perplexity of 15.66, substantially higher than other methods. Table A5 in Appendix B.4 showcases visualization examples, further demonstrating the enhanced performance of SOUL.
@@
BLEU Rouge-L BLEU Rouge-L
@@
Table A7 compares the performance of SOUL with its FO counterparts in the model detoxification task. Similar conclusions can be drawn for both LLaMA2-7B and smaller models such as OPT-350M, consistent with findings from the TOFU and copyright removal tasks.
5.4 Iterative unlearning benefits from SOUL
We next explain the advantage of SOUL over FO optimization-based unlearning methods (such as GA and GradDiff) by examining unlearning and retaining convergence against optimization epochs. Figure 2 shows the forget accuracy (lower values indicate better unlearning efficacy consistent as shown in Table. 2) and retain accuracy (higher values indicate better utility) against the epoch number in the TOFU unlearning task. As we can see, both GA and GradDiff exhibit slower unlearning convergence compared to SOUL (implemented by SO-GradDiff in Table 2). GradDiff, while better at preserving retain accuracy, still falls short in unlearning performance. In contrast, SOUL quickly achieves better forget performance and adaptively adjusts retaining performance, unlike GA, which causes a significant drop in retention at the last epoch. The benefit of SOUL lies in its fast unlearning convergence by accounting for the impact of forget data in (11) and its ability to rewind retaining performance through the adaptive learning rate provided by the second-order optimizer.
Figure 2: Unlearning performance versus optimization epochs using different optimizers in TOFU unlearning. Left: forget accuracy vs. epochs; Right: retain accuracy vs. epochs.
To further justify the iterative unlearning benefit of SOUL, Table A8 compares it with the traditional influence unlearning (IU) method on TOFU. This comparison shows that static IU fails to achieve satisfactory effectiveness due to its lack of optimization power. In contrast, SOUL improves IU by transitioning to an iterative, optimization-driven approach. Additionally, Table A9 shows that SOUL exhibits better unlearning robustness than FO methods in the presence of jailbreak prompts obtained following Lynch et al. (2024). Further, Table A10 presents the time cost of SOUL, demonstrating that the obtained benefits do not come at a substantial cost in time efficiency. This efficiency is due to Sophia leveraging an efficient Hessian diagonal estimate, which avoids the extensive computation typically required for second-order optimization.
6 Conclusion
@@ 8 Acknowledgement
Appendix A Algorithm
Algorithm 1 SOUL to solve problem (4)
2:for $t=1$ to $T$ do
7:     Based on $\mathbf{m}_{t}$ and $\mathbf{h}_{t}$, update $\bm{\theta}$ based on (13):
$\bm{\theta}_{t}=\left\{\begin{array}[]{r}\bm{\theta}_{t-1}+\eta_{t}\mathrm{clip}(\mathbf{m}_{t}/\mathrm{max}\left\{{\gamma\mathbf{h}_{t}},\epsilon\right\},1)\\ \text{(ascent mode for forget data)}\\ \bm{\theta}_{t-1}-\eta_{t}\mathrm{clip}(\mathbf{m}_{t}/\mathrm{max}\left\{{\gamma\mathbf{h}_{t}},\epsilon\right\},1)\\ \text{(descent mode for retain data)}\end{array}\right.$ (A5)
8:end for
When considering PO-type problems like (6), step 7 of Algorithm 1, as depicted in (A5), can only operate in the descent mode. This is because the preference (i.e., the unlearning response $y_{\mathrm{f}}$) has already been defined, and the corresponding forget loss is minimized rather than maximized in (4). In this scenario, SOUL enables the optimization of both forget loss and retain loss through descent mode unification.
Appendix B Additional Experimental Details and Results
B.1 Datasets, tasks and models
Our experimentation revolves around three well-established LLM unlearning tasks. (1) TOFU: This task focuses on fictitious unlearning Maini et al. (2024), involving a dataset of fictitious author profiles for finetuning, and a subset of these profiles constitutes the forget set. We form a forget set by selecting a 10% forget ratio, which includes 400 examples providing information about 20 authors, along with the remaining data points to form the retain set. (2) Copyrighted information removal: This task evaluates the effectiveness of unlearning methods in reducing potential copyright infringement Eldan and Russinovich (2023). We extract 200 chunks from the Harry Potter book series dataset Eldan and Russinovich (2023), with each chunk containing up to 512 tokens, to create the forget set. (3) Model detoxification: This task aims to prevent LLMs from generating toxic content Yao et al. (2023); Ilharco et al. (2022); Zhang et al. (2023c) by employing unlearning approaches. We include 200 negative samples from the PKU-SafeRLHF training set Ji et al. (2024) as the forget set. The C4 dataset Raffel et al. (2020) is used as the retain set for copyright removal and model detoxification tasks to ensure the preservation of model utility.
We selected the OPT-1.3B Zhang et al. (2022a) and LLaMA2-7b Touvron et al. (2023) as foundational models for our study. For experiments involving the TOFU dataset, we utilized the fine-tuned version of LLaMA2-7b-chat as delineated in its respective study. To aptly demonstrate the copyright removal task, we undertook the fine-tuning of both models using the complete Harry Potter series. The fine-tuning procedure for the OPT-1.3B model involved a learning rate of $5\times 10^{-5}$ and a batch size of 2. Conversely, for LLaMA2-7b, we applied Low-Rank Adaptation (LoRA) fine-tuning with a learning rate of $1\times 10^{-4}$ and the same batch size. AdamW served as the optimizer for preparing these models. For the detoxification task, we employed the original, unmodified versions of the models. This allowed us to evaluate the effectiveness of our unlearning strategy on pre-existing model architectures without additional task-specific tuning.
B.2 Unlearning configurations.
LLM unlearning methods and implementation details.
We will assess the effectiveness of our proposed second-order unlearning approach by comparing it with a series of state-of-the-art (SOTA) LLM unlearning techniques. As illustrated in Sec. 3, we consider GradDiff, PO, and NPO, executed via regularized optimization and employing either FO (first-order) optimization or SOUL. We also consider Gradient ascent (GA), which serves as a specialization of GradDiff (4) by setting its regularization parameter $\lambda=0$. In the implementation of PO, we choose a reject-based answer as the target response $y_{\mathrm{f}}$ to steer the model away from unwanted responses. Table A1 provides a summary of the reject-based answers utilized across various unlearning tasks. In addition to the aforementioned finetuning-based unlearning methods, we also explore an input prompt-enabled unlearning approach proposed by Thaker et al. (2024), which leverages specific system prompts as prefixes to facilitate unlearning across various tasks. Further details on these system prompts are provided in Table A2. AdamW Loshchilov and Hutter (2017) is used as the FO optimizer, and Sophia Liu et al. (2023a) (with the default hyperparameter settings) is utilized as the SO optimizer in our proposed SOUL framework presented in Algorithm 1. Table A1 shows the reject-based answers we designed in the preference optimization method: